# May 24, 2024

### v1

@@ 1 Introduction
By construction, the usual PAs are not able to explore the Minkowski region coming from the Euclidean domain as they contain only poles and zeros, and the branch cuts can be, at best, emulated by the accumulation of singularities [44, 45, 46]. Therefore, we propose to accompany the PAs study with the use of a variant of the method called D-Log Padé approximants, D-Logs [32, 47, 48], an extension of PAs which contain, by construction, not only poles and zeros but also branch cuts. A systematic way to bound the value of the function is also provided pointing towards a similar convergence theorem. We sketch a proof of this conjuecture in Appendix B. This can pave the way for future explorations where one could have a glimpse of the Minkowski region from fits in the Euclidean region in a model-independent and systematic way.
@@
This paper is organized as follows. In Sec. 2, we introduce the basic elements to describe the HVP contribution to the $(g-2)_{\mu}$ related to the running of the electromagnetic coupling constant. In Sec. 3, we review the aspects of Padé Theory that are the foundations of our work, in particular, Stieltjes functions and the convergence theorems of PAs and D-Logs applied to them. In Sec. 4, we present the model of Ref. [40] that we use to generate our toy data sets for the Euclidean HVP, while an example of the power of the convergence theorems is presented in Sec. 5, where PAs and D-Logs are built from the exactly known Taylor series given by the model of Sec. 4. In Sec. 6, we employ our method in the idealized scenario where the data points have zero error. Sec. 7 illustrates the application of our method to realistic data sets, following the expectations of the MUonE experiment. Our conclusions are given in Sec. 8. Technical details about the fitting functions and convergence theorems are relegated to the Appendices A and B, correspondingly.
@@ A PA $P_{M}^{N}(z)$ is a rational function given by
are positive for any $m\geq 0$ and $n\geq 0$. These determinants produce constraints between the Taylor series coefficients $f_{n}$ of the Stieltjes functions. They will also prevent the appearance of defects, or technically Froissart doublets, exact cancellations between the zeros of the numerator and the denominator [44], effectively reducing the order of the approximants. These will be essential in our analysis since a PA or D-Log constructed from a Stieltjes function has the same Stieltjes properties as the original function has [43, 44]. The D-Log in Eq. (14) is, in general, not a rational approximant, but the function $F(z)$ is meromorphic and, because of that, it is easily approximated by the PA $\bar{P}_{M}^{N}$, and the convergence of the approximation can be proven [43]. What is more, if $f(z)$ is a Stieltjes function, we conjecture that $F(z)$ will also be Stieltjes, thus convergence theorems for Stieltjes functions apply to the approximation of $F(z)$ by $\bar{P}_{M}^{N}$ as well. A sketch of a proof of this conjecture is given in Appendix B. The exception to the convergence rule is the $P^{N}_{1}$ sequence as one can prove that approximants belonging to this sequence are not necessarily Stieltjes even when built to approximate a Stieltjes function [44].
@@
For the D-Log case, a similar pattern is found. The sequence $D_{N}^{N}$ together with $D^{N}_{N+1}$ and $D_{N}^{N+1}$ bound the original function. In our case of study, we found the fastest convergence with the $D^{N}_{N+1}$ sequence, which renders the theorem as
@@ 4 A model for the Euclidean correlator
We consider this simple model to be sufficiently realistic because it leads to a representation of $\Delta\alpha_{\mathrm{had}}$ which is a Stieltjes function, as expected in QCD. Hence, the theorems given in Sec. 3 and the coefficient constraints of Eq. (21) are all valid.
@@ 5.2 Convergence of D-Logs and comparison with PAs results
We can also build D-Logs to the Taylor series of $\Delta\alpha_{\mathrm{had}}(t)$. As stated by the convergence theorems, specific sequences of D-Logs to $\Delta\alpha_{\mathrm{had}}(t)$ bound the function from above and from below, as can be seen in Fig. 4, in complete analogy to the convergence pattern of PAs shown in Figs. 2a and 2b.
@@ Acknowledgements
The authors thank Carlo Carloni Calame, Giovanni Abbiendi and Graziano Venanzoni for providing us with the data errors expected in the MUonE experiment. PM thanks IFSC, at University of Sao Paulo, and CYL thanks IFAE, at Universitat Autònoma de Barcelona, for hospitality. The work of CYL was financed by the São Paulo Research Foundation (FAPESP) grant No. 2020/15532-1 and 2022/02328-2. DB’s work was supported by FAPESP grant No. 2021/06756-6 and by CNPq grant No. 308979/2021-4. The work of PM has been supported by the Ministerio de Ciencia e Innovación under grant PID2020-112965GB-I00 and by the Secretaria d’Universitats i Recerca del Departament d’Empresa i Coneixement de la Generalitat de Catalunya under grant 2021 SGR 00649. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.
@@ Appendix A Fitting functions constructed using PAs and D-Logs
Appendix B Convergence of D-Log approximants
Little is known about the convergence of D-Log approximants to a given function. In this Appendix, we show that D-Log convergence to Stieljtes functions can be obtained using the same convergence theorems for PAs to Stieltjes functions. The rationale is the following. Since D-Logs are built from the exponentiation and integration over a $\bar{P}_{M}^{N}(z)$ to $F(z)=\frac{{\rm d}}{{\rm d}z}\ln f(z)$, if $F(z)$ is a Stieltjes function, then the sequences $\bar{P}_{N}^{N}(z)$ and $\bar{P}^{N\pm 1}_{N}(z)$ will bound $F(z)$. Integrating and exponentiating do not disturb the hierarchy of the sequences yielding the bounds for the D-Logs exploited in this work.
The question to answer now is whether the logarithmic derivative of a Stieltjes function is a Stieltjes function as well. A necessary and sufficient condition for a function $f(z)$ to be Stieltjes is that all determinants $D_{m,n}^{(f)}$, made out of the moments $f_{n}$, are positive. Then, to prove that $F(z)$ is a Stieljtes function, we need to verify that its determinants $\mathscr{D}_{m,n}^{(F)}$ are all positive. Since they are not known, this may seem an impossible task. We know, nonetheless, that the determinants $\mathscr{D}_{m,n}^{(f)}$ of $f(z)$ are all positive. We can then relate the determinants of $F(z)$ with the ones of $f(z)$ and use the positivity condition of the latter to unravel the positivity condition of the former.
The definition of a Stieltjes function gives $f_{n}=\int_{0}^{\infty}u^{n}\mathrm{d}\phi(u)$, c.f. Eq. (16). The distance between consecutive coefficients $f_{n}$ and $f_{n+1}$ can only be a function of $n$ (irrespective whether the distance grows or decreases):
$f_{n}-f_{n+1}=\int_{0}^{\infty}u^{n}(1-u)\mathrm{d}\phi(u)\,.$ (44)
This implies that either $f_{n}>f_{n+1}$ or $f_{n}<f_{n+1}$. The determinants of $f(z)$ are informative. In particular, the $\mathscr{D}_{n,0}^{(f)}>0$ imply all $f_{n}>0$ and Eq. (44) implies $D_{n,1}^{(f)}=f_{n}f_{n+2}-f_{n+1}^{2}>0$. Similarly, combinations such as $f_{n}f_{n+3}-f_{n+1}f_{n+2}$ or $f_{0}f_{n-1}-f_{1}f_{n-2}$ are positive for all $n$. The reason is the following. We focus on the case where $f_{n}>f_{n+1}$ and, as stated, the distance between two consecutive $f_{n}$ is $\sim 1/n$. The product $f_{n}f_{n+3}-f_{n+1}f_{n+2}$ represents the difference between the surface of two rectangles. Comparing sizes, $f_{n}$ is larger than $f_{n+1}$ by $1/n$ while $f_{n+2}-f_{n+3}\sim 1/(n+2)$. The surface of the first rectangle is then larger and the difference of surfaces is positive. If $f_{n}$ coefficients grow with $n$, the argument is essentially the same.
We show now how the determinants of $F(z)$ can be built from the determinants of $f(z)$. Notice that $F(z)$ is the logarithm derivative of $f(z)$ and then
$F(z)=\frac{f^{\prime}(z)}{f(z)}\,=\sum_{n=0}^{\infty}F_{n}z^{n}=\frac{\sum_{n=1}^{\infty}n\,f_{n}\,z^{n-1}}{\sum_{n=0}^{\infty}\,f_{n}\,z^{n}}\,.$ (45)
The coefficients $F_{n}$ are built from the Taylor coefficients of the product of two functions: $f^{\prime}(z)$ and $1/f(z)$. The Taylor expansion around $z=0$ of the product of two generic analytic functions $g(z)$ and $h(z)$ can be written in the compact form:
$g(z)h(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{n!}\sum_{i=0}^{n}\binom{n}{i}g_{i}\,h_{n-i}\,.$ (46)
This expansion shows that the $n$-th coefficient of the product of two functions contains $n+1$ terms. A closer look at them reveals a hierarchy:
$F_{n}=\sum_{i=0}^{n}(n+1-i)f_{n+1-i}\left(\frac{1}{f(z)}\right)_{i},$ (47)
where $\left(\frac{1}{f(z)}\right)_{i}$ refers to the $i$-th Taylor coefficient of $1/f(z)$. The $i=0$ term of $1/f(z)$ is simply $1/f_{0}$.
The terms in $F_{n}$ from Eq. (47) have alternate sign as the $\left(\frac{1}{f(z)}\right)_{i}$ contributes with $(-1)^{i}$ (since all $f_{n}>0$). One can group them in terms of determinants of $f(z)$ for which the sign is known. The value of the coefficient $F_{n}$ is dominated by the combination $\left((n+1)(f_{0}f_{n+1}-f_{n}f_{1})f_{0}^{n-1}+f_{1}^{n+1}\right)/f_{0}^{(n+1)}$ which does not belong to any determinant and is positive. This particular combination arises from the $i=0$ and $i=n$ terms defining $F_{n}$. The rest of $i=1,2,...,n-1$ coefficients can be factorized in terms of $\mathscr{D}_{m,1}^{(f)}$. Since the determinants are very small, numerically, compared to coefficients $f_{n}$, terms proportional to a determinant are subleading. Since the dominant contribution is positive, $F_{n}>0\,\forall n$. $F_{n}$ coefficients from an exemplary $f(z)$ function are given in the next subsection.
For the calculation of the determinants $\mathscr{D}_{n,m}^{(F)}$ of $F(z)$, we keep the dominant terms. The $i=0$ coefficient of $F_{n}$ is proportional to $(n+1)f_{n+1}$. This coefficient dominates as it is the biggest one, and positive, which ensures the determinants $\mathscr{D}_{n,1}^{(F)}$ are positive (using arguments similar to $\mathscr{D}_{n+1}^{(f)}$). In detail, the dominant term in $\mathscr{D}_{n,1}^{(F)}$ reads
$\mathscr{D}_{n,1}^{(F)}=$ $\frac{(n+1)(f_{n+1}f_{0}-f_{n}f_{1})}{f_{0}^{2}}\frac{(n+3)(f_{n+3}f_{0}-f_{n+2}f_{1})}{f_{0}^{2}}$
$-\frac{(n+2)^{2}(f_{n+2}f_{0}-f_{n+1}f_{1})^{2}}{f_{0}^{4}}+\cdots.$ (48)
Since $(n+1)(n+3)=(n+2)^{2}-1$, we can argue $(n+1)(n+3)\sim(n+2)^{2}$, group the two terms of the $\mathscr{D}_{n,1}(F)$, and factorize the terms proportional to $f_{0}^{2},f_{1}^{2}$ and $f_{0}f_{1}$:
$\mathscr{D}_{n,1}(F)=$ (49)
$=\frac{(n+2)^{2}}{f_{0}^{4}}\left(f_{0}^{2}(f_{n+1}f_{n+3}-f_{n+2}^{2})+f_{1}^{2}(f_{n}f_{n+2}-f_{n+1}^{2})-f_{0}f_{1}(f_{n+3}f_{n}-f_{n+1}f_{n+2})\right).$
The coefficient of the term proportional to $f_{0}^{2}$ is the $\mathscr{D}_{n+1,1}^{(f)}$ of $f(z)$ and that of the term proportional to $f_{1}^{2}$ is the $\mathscr{D}_{n,1}^{(f)}$. The coefficient of the term proportional to $f_{0}f_{1}$ is not a determinant. Using the hierarchy of coefficients deduced from Eq. (44) we observe it is positive but smaller than the other two terms. We conclude $\mathscr{D}_{n,1}^{(F)}>0,\forall n$.
The next step is related to the positivity condition that the higher-order determinants $\mathscr{D}_{n,m}^{(F)}$ satisfy. By Cramer’s rule, $\mathscr{D}_{n,m}^{(F)}$ can be calculated by a sign-alternating sum of different contributions as happens with the construction of the $\mathscr{D}_{n,1}(F)$ in Eq. (49). Among them, few are $\mathscr{D}_{n,m-1}^{(F)}$ determinants multiplied by the corresponding $F_{n}$ coefficient, while the others are combinations of coefficients which are numerically smaller than $\mathscr{D}_{n,m-1}^{(F)}$. When applying Cramer’s rule, one can always start with the term that contains the coefficient $F_{n}$ with the smallest $n$. The subsequent terms compete and are sign alternating. The term with smallest $n$ dominates and defines the sign of $\mathscr{D}_{n,m}^{(F)}$. As an example, $\mathscr{D}_{0,2}^{(F)}$ reads
$\mathscr{D}_{0,2}^{(F)}$ $=F_{0}(F_{2}F_{4}-F_{3}^{2})-F_{1}(F_{1}F_{4}-F_{2}F_{3})+F_{2}(F_{1}F_{3}-F_{2}^{2})$
$=F_{0}\mathscr{D}_{2,1}^{(F)}-F_{1}(F_{1}F_{4}-F_{2}F_{3})+F_{2}\mathscr{D}_{1,1}^{(F)}>0.$ (50)
Since $\mathscr{D}_{n,1}^{(F)}>0$, $\mathscr{D}_{n,1}^{(F)}>F_{n}F_{n+3}-F_{n+1}F_{n+2}>0$ by virtue of Eq. (44) as argued before, and $F_{n}>F_{n+1}$, $\mathscr{D}_{n,m}^{(F)}>0$, for all $n$.
B.1 Example
We illustrate the proof sketched above with an example. Let $f(z)=\frac{\log[1+z]}{z}$ for which $\phi(u)=u$ in $u\in[0,1]$ and $\phi(u)=0$ in $u\in(1,\infty)$. As such, $f_{n}=\int_{0}^{1}u^{n}\mathrm{d}\phi(u)=\frac{1}{n+1}$. In this case:
  • $\mathscr{D}_{n,0}^{(f)}>0$ as $f_{n}=\frac{1}{n+1}>0$ $\forall n$
  • $\mathscr{D}_{n,1}^{(f)}=f_{n}f_{n+2}-f_{n+1}^{2}=\frac{1}{(n+1)(n+2)^{2}(n+3)}\sim\frac{1}{(n+2)^{4}}>0$ $\forall n$. Notice $\mathscr{D}_{n,0}^{(f)}\gg\mathscr{D}_{n,1}^{(f)}$.
  • $f_{n}-f_{n+1}=\int_{0}^{1}u^{n}(1-u){\rm d}u=\frac{1}{n^{2}+3n+2}>0$ $\forall n$.
  • $f^{\prime}_{n}=\frac{n}{n+1}>0$ $\forall n$.
  • $(1/f)_{n}<0$ $\forall n$.
With this at hand, the Taylor coefficients $F_{n}$ of the function $F(z)=\frac{f^{\prime}(z)}{f(z)}$ read
  • $F_{0}=\frac{f_{1}}{f_{0}}=\frac{1}{2}$,
  • $F_{1}=\frac{2f_{0}f_{2}-f_{1}^{2}}{f_{0}^{2}}=\frac{f_{0}f_{2}+(f_{0}f_{2}-f_{1}^{2})}{f_{0}^{2}}=\frac{1}{3}+\left(\frac{1}{3}-\frac{1}{2^{2}}\right)=\frac{1}{3}+\frac{1}{12}=\frac{5}{12}$,
  • $F_{2}=\frac{3(f_{3}f_{0}-f_{2}f_{1})f_{0}+f_{1}^{3}}{f_{0}^{3}}=3\left(\frac{1}{4}-\frac{1}{3}\,\frac{1}{2}\right)+\frac{1}{2^{3}}=\frac{2}{8}+\frac{1}{8}=\frac{3}{8}$,
  • $F_{3}=\frac{4(f_{4}f_{0}-f_{3}f_{1})f_{0}^{2}+2f_{0}f_{2}(f_{1}^{2}-f_{0}f_{2})-f_{1}(f_{1}^{3}-2f_{0}f_{1}f_{2})}{f_{0}^{4}}=\frac{4(f_{4}f_{0}-f_{3}f_{1})f_{0}^{2}-2(f_{1}^{2}-f_{0}f_{2})^{2}+f_{1}^{4}}{f_{0}^{4}}=\frac{3}{10}-\frac{2}{144}+\frac{1}{16}=\frac{251}{720}$,
  • $F_{4}=\frac{5(f_{5}f_{0}-f_{4}f_{1})f_{0}^{3}+5f_{3}f_{0}^{2}(f_{1}^{2}-f_{0}f_{2})-5f_{2}f_{1}f_{0}(f_{1}^{2}-f_{0}f_{2})+f_{1}^{5}}{f_{0}^{5}}=\\ \frac{5(f_{5}f_{0}-f_{4}f_{1})f_{0}^{3}+5f_{0}(f_{3}f_{0}-f_{2}f_{1})(f_{1}^{2}-f_{0}f_{2})+f_{1}^{5}}{f_{0}^{5}}=\frac{1}{3}-\frac{5}{144}+\frac{1}{32}=\frac{95}{288}$,
  • $F_{5}=\frac{6(f_{6}f_{0}-f_{5}f_{1})f_{0}^{4}+[6f_{0}^{2}(f_{4}f_{0}-f_{3}f_{1})-6f_{0}f_{1}(f_{3}f_{0}-f_{2}f_{1})](f_{1}^{2}-f_{0}f_{2})}{f_{0}^{6}}+\\ \frac{-2(f_{1}^{2}-f_{0}f_{2})^{3}+3f_{0}(f_{3}f_{0}-f_{2}f_{1})(f_{1}^{3}+f_{1}(f_{1}^{2}-f_{0}f_{2})-f_{0}^{2}f_{3})+f_{1}^{6}}{f_{0}^{6}}=\frac{19087}{60480}$.
The recurrence pattern is clear: each $F_{n}$ ($n>2$) has a term $(n+1)(f_{0}f_{n+1}-f_{1}f_{n})f_{0}^{n-1}=\frac{n}{2(n+2)}>0$ that dominates over the negative terms which, in their turn, are proportional to $\mathscr{D}_{n,1}^{(f)}$. Since $\mathscr{D}_{n,1}^{(f)}\sim\frac{1}{(n+2)^{4}}\ll 1$, all the terms proportional to the determinants appearing in the construction of the $F_{n}$ coefficients are negligible.
For the $\mathscr{D}_{n,1}^{(F)}$ determinants built for $F(z)$, we exploit the same hierarchy. Take the $F_{n}$ coefficients, neglect the terms within them proportional to the determinants from $f(z)$ and also the one proportional to $f_{1}^{n+1}$. Then immediately one recognizes that $\mathscr{D}_{n,1}^{(F)}>0$ $\forall n$. For example, $\mathscr{D}_{n=1,1}^{(F)}=F_{1}F_{3}-F_{2}^{2}\sim\frac{1}{3}\frac{n+2}{2(n+4)}-\left(\frac{n+1}{2(n+3)}\right)^{2}\big{|}_{n=1}=\frac{3}{80}$. Similarly, $\mathscr{D}_{n=2,1}^{(F)}=F_{2}F_{4}-F_{3}^{2}\sim\frac{3}{8}\frac{95}{288}-\left(\frac{251}{720}\right)^{2}=\frac{281}{129600}\sim\frac{3}{8}\frac{1}{3}-\left(\frac{3}{10}\right)^{2}=\frac{7}{200}$. We can explicitly calculate Eq. (49) for arbitrary $n$. Take now $n=3$:
$\mathscr{D}_{3,1}^{(F)}=\frac{(3+2)^{2}}{1}\left[1^{2}\left(\frac{1}{5}\,\frac{1}{7}-\frac{1}{6^{2}}\right)+\frac{1}{2^{2}}\left(\frac{1}{4}\,\frac{1}{6}-\frac{1}{5^{2}}\right)+1\,\frac{1}{2}\left(\frac{1}{5}\,\frac{1}{6}-\frac{1}{7}\,\frac{1}{4}\right)\right]=\frac{1}{2016}$ (51)
For $n>4$, the dominant terms for each $F_{n}$ allow a compact expression for the determinant: $\mathscr{D}_{n,1}^{(F)}=F_{n}F_{n+2}-F_{n+1}^{2}=\frac{n-4}{4(n+3)(n+4)}>0$.
Higher-order determinants $\mathscr{D}_{m,n}^{(F)}$ can be constructed following the results from the lowest ones using the aforementioned Cramer’s rule method, thus returning them all positive. As an example, $\mathscr{D}_{0,2}^{(F)}$ reads
$\mathscr{D}_{0,2}^{(F)}=F_{0}\mathscr{D}_{2,1}^{(F)}-F_{1}(F_{1}F_{4}-F_{2}F_{3})+F_{2}\mathscr{D}_{1,1}^{(F)}=\frac{7}{400}-\frac{29}{10368}+\frac{9}{640}>0\,.$ (52)
We can conclude that $F(z)$ is a Stieltjes function since all its determinants are positive, as we wanted to demonstrate.

### v2

# October 15, 2024

@@ 1 Introduction
By construction, the usual PAs are not able to explore the Minkowski region coming from the Euclidean domain as they contain only poles and zeros, and the branch cuts can be, at best, emulated by the accumulation of singularities [44, 45, 46]. Therefore, we propose to accompany the PAs study with the use of a variant of the method called D-Log Padé approximants, D-Logs [32, 47, 48], an extension of PAs which contain, by construction, not only poles and zeros but also branch cuts. In some cases, a systematic way to bound the value of the function is also provided pointing towards a similar convergence theorem. This can pave the way for future explorations where one could have a glimpse of the Minkowski region from fits in the Euclidean region in a model-independent and systematic way.
@@
This paper is organized as follows. In Sec. 2, we introduce the basic elements to describe the HVP contribution to the $(g-2)_{\mu}$ related to the running of the electromagnetic coupling constant. In Sec. 3, we review the aspects of Padé Theory that are the foundations of our work, in particular, Stieltjes functions and the convergence theorems of PAs and D-Logs applied to them. In Sec. 4, we present the model of Ref. [40] that we use to generate our toy data sets for the Euclidean HVP, while an example of the power of the convergence theorems is presented in Sec. 5, where PAs and D-Logs are built from the exactly known Taylor series given by the model of Sec. 4. In Sec. 6, we employ our method in the idealized scenario where the data points have zero error. Sec. 7 illustrates the application of our method to realistic data sets, following the expectations of the MUonE experiment. Our conclusions are given in Sec. 8. Technical details about the fitting functions are relegated to Appendix A.
@@ A PA $P_{M}^{N}(z)$ is a rational function given by
are positive for any $m\geq 0$ and $n\geq 0$. These determinants produce constraints between the Taylor series coefficients $f_{n}$ of the Stieltjes functions. They will also prevent the appearance of defects, or technically Froissart doublets, exact cancellations between the zeros of the numerator and the denominator [44], effectively reducing the order of the approximants. These will be essential in our analysis since a PA or D-Log constructed from a Stieltjes function has the same Stieltjes properties as the original function has [43, 44]. The D-Log in Eq. (14) is, in general, not a rational approximant, but the function $F(z)$ is meromorphic and, because of that, it is easily approximated by the PA $\bar{P}_{M}^{N}$, and the convergence of the approximation can be proven [43]. What is more, for some Stieltjes functions $f(z)$, it can be shown that $F(z)$ will also be Stieltjes. In these cases, convergence theorems for Stieltjes functions apply to the approximation of $F(z)$ by $\bar{P}_{M}^{N}$ as well. The exception to the convergence rule is the $P^{N}_{1}$ sequence as one can prove that approximants belonging to this sequence are not necessarily Stieltjes even when built to approximate a Stieltjes function [44].
@@
For the D-Log case, a similar pattern is found, in our case. The sequence $D_{N}^{N}$ together with $D^{N}_{N+1}$ and $D_{N}^{N+1}$ bound the original function. We found the fastest convergence with the $D^{N}_{N+1}$ sequence, following the pattern
@@ 4 A model for the Euclidean correlator
We consider this simple model to be sufficiently realistic because it leads to a representation of $\Delta\alpha_{\mathrm{had}}$ which is a Stieltjes function, as expected in QCD. Hence, the theorems given in Sec. 3 and the coefficient constraints of Eq. (21) are all valid. We also note that numerical checks of several hundred determinants of Eq. (17) built from the series of $\frac{d}{dt}\ln\Delta\alpha_{\rm had}(t)$ indicate that this function, which is required to construct the D-Logs, is also Stieltjes.
@@ 5.2 Convergence of D-Logs and comparison with PAs results
We can also build D-Logs to the Taylor series of $\Delta\alpha_{\mathrm{had}}(t)$. As stated by the convergence theorems for Stieltjes functions, specific sequences of D-Logs to $\Delta\alpha_{\mathrm{had}}(t)$ bound the function from above and from below, as can be seen in Fig. 4, in complete analogy to the convergence pattern of PAs shown in Figs. 2a and 2b.
@@ Acknowledgements
The authors thank Carlo Carloni Calame, Giovanni Abbiendi and Graziano Venanzoni for providing us with the data errors expected in the MUonE experiment. We thank Jérôme Charles for comments on a previous version of this manuscript. PM thanks IFSC, at University of Sao Paulo, and CYL thanks IFAE, at Universitat Autònoma de Barcelona, for hospitality. The work of CYL was financed by the São Paulo Research Foundation (FAPESP) grant No. 2020/15532-1 and 2022/02328-2. DB’s work was supported by FAPESP grant No. 2021/06756-6 and by CNPq grant No. 308979/2021-4. The work of PM has been supported by the Ministerio de Ciencia e Innovación under grant PID2020-112965GB-I00 and by the Secretaria d’Universitats i Recerca del Departament d’Empresa i Coneixement de la Generalitat de Catalunya under grant 2021 SGR 00649. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.